What force is required to move a load of 5kg through a distance of 1000cm in 5seconds
1
Expert's answer
2020-02-26T10:09:35-0500
"F=ma"

"a=\frac{2s}{t^2}"

"F=m\frac{2s}{t^2}=5\frac{2(10)}{5^2}=4\ N"